OpenStudy (anonymous):

Solve for x: (x+20)/2 = 3x x = 10 x = 4 x = -14 x = -17

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{x+20}{2}=3x }\) \(\LARGE\color{blue}{ \frac{x+20}{2}=\frac{6x}{2} }\) \(\Large\color{blue}{ x+20=6x }\) \(\Large\color{blue}{ 20=5x }\) \(\Large\color{blue}{ x=? }\)

OpenStudy (solomonzelman):

let me know if you need further assistance.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x=4 ?

OpenStudy (solomonzelman):

Yes, x=4 is correct !

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

Anytime!